# The GECKOS Survey: Identifying kinematic sub-structures in edge-on galaxies

### v1

@@ 3.4.3 $h_{3}$ maps
The central $h_{3}$ feature detected in the 2D maps is also seen in the 1D radial profiles of Figure 16a. A strong, negative gradient in $h_{3}$ as a function of galactic radius is seen in the $h_{3}$ profiles of PGC 044931, NGC 3957, IC 1711, IC 5244, and UGC 00903.
3.4.4 $h_{4}$ maps
In Figures 314, we present the line-of-sight kurtosis, or $h_{4}$, maps. There are fewer spatially coherent features visible in these maps, possibly the result of observational challenges related to retrieving high-order velocity moments from medium-resolution spectra (e.g. Wang et al., 2024). Nevertheless, we do notice some trends. We expect that the large longitudinally-extended regions of high positive $h_{4}$ around the centres of some galaxies and corresponding to low $\sigma_{\star}$ are caused by the dust lanes. In some galaxies, including PGC 044931, NGC 3957, NGC 0522, and ESO 079-003, we see regions of high positive $h_{4}$ extending from the central regions of the galaxy 1–2 kpc off the midplane. In PGC 044931, IC 1711, and NGC 3957, the small, inner areas of $h_{3}$-$V_{\star}$ anti-correlation is co-spatial with a coherent region of high positive $h_{4}$. The strength of these features with respect the the surrounding disc are highlighted in the 1D profiles of Figure 16b.
@@ 4 Discussion
Here we attempt to connect the observations of Section 3.3 to an understanding of the central kinematic sub-structures in the GECKOS galaxies. Already clear is the diversity of structures in the $V_{\star}$ maps (visible central kinematically-decoupled rapidly-rotating structures), $\sigma_{\star}$ maps (structure and intensity of sigma peaks at the centre of the galaxy, effects of dust lanes, central ‘croissant’ features), $h_{3}$ (coherent areas of (anti-)correlation with $V_{\star}$), and $h_{4}$ (small, highly-peaked central regions). That said, some features are similar throughout the sample: all galaxies possess a rotating disc structure and some degree of $h_{3}$-$V_{\star}$ anti-correlation in the outer disc regions of the galaxy. In the following, we explore these observations as possible tracers of kinematic sub-structures.
@@ 4.1.1 Midplane features
  • A ‘double-hump’ rotation curve, with a stellar mean line-of-sight velocity ($V_{\star}$) profile exhibiting an initial rise, followed by a hump or plateau, and then a secondary rise, before flattening out. The hump is caused by the bar $x_{1}$ orbits, the elongated family of orbits that follow the major axis of the bar (e.g. Contopoulos & Papayannopoulos, 1980; Athanassoula, 1992), adding a distinct component to the LOSVD.
  • A flat-topped or weakly peaked radial velocity dispersion ($\sigma_{\star}$) profile, sometimes with a local central minimum. Broad ‘shoulders’ or secondary maxima may be present in strong bars, usually towards the end of the bar, just inside the inner ring. Bureau & Athanassoula (2005) attributed the central $\sigma_{\star}$ peak to the large variety of orbital shapes encountered in strongly-barred discs, particularly along the major axis. They posited that the $\sigma_{\star}$ ‘shoulders’ are produced by the tips of the $x_{1}$ orbits and the inner 4:1 orbits. In particular, $x_{1}$ orbits with loops around the major axis are expected in strong bars, causing a local increase in the velocity dispersion, strongest for bars seen side-on (Athanassoula, 1992). The higher energy inner 4:1 orbits have similar loops, which increase $\sigma_{\star}$ for bars seen both side-on and end-on.
@@
Crucially, Bureau & Athanassoula (2005) also show that the strength of these kinematic features is correlated with bar viewing angle, such that the strongest signals are seen when the bar is oriented end-on to the viewer.
@@
Comparing the $h_{3}$ 1D profiles of Figure 16a with $V_{\star}$ in Figure 15a, some $h_{3}-V_{\star}$ correlations are seen in the central few kpc of PGC 044931, NGC 3957, IC 1711, IC 5244, ESO 484-036, and ESO 079-003. Taken with the trends seen in the $V_{\star}$ and $\sigma_{\star}$ profiles, we can reliably report that these galaxies host strong evidence of bar structure from their midplane stellar kinematics.
@@ 4.1.2 Off-plane features
Iannuzzi & Athanassoula (2015) extended previous long-slit work by providing 2D, projected line-of-sight kinematics for a set of barred/BP bulge dynamical models featuring star formation, one of which also hosted a spherical bulge. In addition to confirming 1D predictions and observations, they found that the 2D line-of-sight mean velocity maps do not provide a robust predictor of the presence of BP bulges as the degree of cylindrical rotation (i.e. mean stellar velocity independent of height above the disc) varies. In their work, Iannuzzi & Athanassoula (2015) reported a tendency for the degree of cylindrical rotation to weaken when simulations are seen from side-on to progressively more end-on bar angles, though the importance of this effect varies from case to case without a clear trend in BP bulge strength. A variation in BP bulge rotational properties is also reported by Athanassoula & Misiriotis (2002), who attribute this behaviour to different disc and halo contributions to the total circular velocity curve in their various models. Williams et al. (2011)’s observations support this idea with observations of five edge-on discs hosting BPs that show a variety of projected off-plane velocity behaviours. However, Molaeinezhad et al. (2016) developed a new technique to measure the degree of cylindrical rotation and found a wide range of cylindrical rotation, but on average, BP bulges displayed higher values for this parameter than non-BP bulges. We do not look further into cylindrical rotation indicators in this work, though note that the GECKOS sample will provide an excellent opportunity to understand the off-plane velocity behaviour of BP bulges.
@@
  • Elongated ‘wings’ of large $h_{3}$. Again, the bar viewing angle is important for both the amplitude and morphology of $h_{3}$ maps. Iannuzzi & Athanassoula (2015) and Li et al. (2018) analyse separate N-body models of barred systems hosting a BP bulge at several different bar viewing angles and conclude that while bar-driven signatures in $h_{3}$ are barely visible when the bar is side-on to the observer, considerable changes occur as the bar is rotated, such that large regions of $h_{3}$ maxima and minima develop roughly in the regions corresponding to the peanut, increasing in size and magnitude as the bar gets closer to end-on. Hence, the largest kinematic signal in $h_{3}$ maps should be seen in end-on bars.
  • ‘X’-shaped region of deep $h_{4}$ minima. This feature was also predicted along the bar major axis in more face-on galaxies by Debattista et al. (2005), and observationally confirmed for one galaxy by Méndez-Abreu et al. (2008) and within an inner bar by Méndez-Abreu et al. (2019). The minima occur because the vertical density distribution of stars becomes broader at the location of the BP bulge, leading to a double minimum of the fourth-order Gauss-Hermite moment of the vertical density distribution. The LOSVD moment $h_{4}$ is then an excellent proxy for the unobservable vertical density distribution moments. This is because the stars composing the $X$-shape sit on top of (in projection) the off-plane stellar distribution, while exhibiting markedly different kinematics. The presence of these two distinct spatially-overlapping kinematic populations produces a ‘flat-topped’ or even double-peaked LOSVD, leading to negative $h_{4}$.
Both $h_{3}$ and $h_{4}$ features become stronger at larger BP bulge strengths in edge-on galaxies. The former are mainly characteristic of intermediate PAs (30-60°), while the latter are visible down to the end-on view.
Taking these predictions to the GECKOS sample, we see coherent areas of $h_{3}$ that correlate with the respective $V_{\star}$ maps at the centres of all BP bulge galaxies (though to varying degrees). We see an excellent example of the predicted wing structure in IC 5244 and IC 1711. These areas of $h_{3}$-$V_{\star}$ correlation originate in the central galactic regions and extend up to 2 kpc off the disc, and up to 5 kpc radially. We expect that these regions correspond to the area of influence of the bar, with the radial width related to the bar viewing angle.
@@
Both Iannuzzi & Athanassoula (2015) applied Voronoi binning to their particle data, before fitting the LOSVD with a Gauss-Hermite to obtain their kinematics; an equivalent strategy to Wang et al. (2024). We believe that the reason that these simulation studies found strong $h_{4}$ features that are not seen in the GECKOS data is most likely due to observational limitations, i.e., the process of transferring particle statistics to spectra, and then extracting the $h_{4}$ from weak spectral features. Due to this process, all of the ‘real’ $h_{4}$ features become more shallow or disappear. We refer the reader to Sections 3.3 and 4.1 of Wang et al. (2024) for a more thorough discussion on this point.
@@ 4.1.3 Bar viewing angle
$h_{3}$ maps vary considerably in both amplitude and morphology with bar viewing angle, according to both the Iannuzzi & Athanassoula (2015) and Li et al. (2018) model predictions. The region of $h_{3}$-$V_{\star}$ correlation appears least conspicuous in the side-on case, and increases in size and magnitude (whilst still maintaining the appearance of extended ‘blobs’), as the bar is rotated toward the viewer. The correlation is most prominent when the bar is seen end-on. In Figures 3-14, we see the strongest degree of coherent $h_{3}$ structure (correlating with $V_{\star}$) in the central regions of IC 5244 and IC 1711, adding weight to the idea that these galaxies possess end-on bars. We see the lowest degree of $h_{3}$ structure in ESO 120-016 and NGC 0522. While it is possible that ESO 120-016 is indeed not a BP bulge galaxy, NGC 0522 possesses a definite X shape in the DECaLS and Spitzer imaging. For this reason, we expect that NGC 0522 must possess a bar that is close to side-on orientation. For all other galaxies, we infer from the $h_{3}$ maps that the bar viewing angle is somewhat intermediate between side-on and end-on.
@@ 4.2 Evidence for nuclear discs
Following from the predictions of Bureau & Athanassoula (2005), an additional predictor of bar structure along the kinematic major axis of edge-on galaxies was provided by Chung & Bureau (2004), who illustrate the features mentioned in Section 4.1 in their Figure 9, and observed a strong anti-correlation between $h_{3}$ and $V_{\star}$ in the innermost region of BP bulge galaxies, which they attributed to the presence of a cold, rapidly-rotating nuclear disc built by a bar. This anti-correlation switches back to a correlation outside of the nuclear disc region but still within the bar region, in line with the Bureau & Athanassoula (2005) prediction.
@@
Evidence for the presence of nuclear discs is observed in the $V_{\star}$, $\sigma_{\star}$, $h_{3}$, and $h_{4}$ maps of IC 1711, IC 5244, NGC 3957, and PGC 044931. Whilst all show modest increases in rotational velocity in small regions at their very centres, NGC 3957 and PGC 044931 also show a ‘croissant’-shaped feature in their velocity dispersion maps, with a central depression of lower $\sigma_{\star}$. All four galaxies also show $h_{3}$-$V_{\star}$ anti-correlations in these central regions, co-located with regions of high positive $h_{4}$.
@@
These features are perhaps better shown in the 1D line profiles of Figures 15 and  16, where the four galaxies in question lie at the bottom of these figures. The central minima in the dispersion profiles of PGC 044931, NGC 3957, and IC 1711 are features that are sometimes referred to as ‘sigma drops’ (e.g. Emsellem et al., 2001). Observationally, this feature has been frequently seen in long-slit spectra (e.g. Emsellem et al., 2001; Márquez et al., 2003; Chung & Bureau, 2004; Comerón et al., 2008; Méndez-Abreu et al., 2014) and IFS data (e.g. Crocker et al., 2010; Lin et al., 2017; Pinna et al., 2019b; Shimizu et al., 2019). We expect that these $\sigma$ drops are due to the presence of a dynamically cold and kinematically-decoupled central stellar disc originating from gas inflow. Interestingly, upon closer inspection of the $\sigma_{\star}$ radial profiles of PGC 044931, NGC 3957, and IC 1711, we see that their profiles bifurcate in the central regions, with one population of stars of low $\sigma_{\star}$ (the nuclear disc stars), and one population with higher $\sigma_{\star}$. In fact, in IC 5244 we see only a $\sigma_{\star}$ peak, with perhaps only a tiny hint of a drop. We expect that the peaks are due to bar orbits, as discussed in Section 4.1.
@@
The N-body models of Bureau & Athanassoula (2005) do not include nuclear discs, but do reproduce this central $\sigma_{\star}$ minima, attributing it to the orbital structure of strongly barred discs. More recent N-body and SPH simulations that include stars, gas, and star formation show that young stars born in the nuclear regions from dynamically cold gas have a velocity dispersion lower than the older stellar population (Wozniak et al., 2003; Michel-Dansac & Wozniak, 2004; Wozniak, 2007; Cole et al., 2014; Portaluri et al., 2017), and hence attribute the sigma drop feature to a dynamically cold, nuclear disc. Contextually, the presence of a nuclear disc in a barred galaxy makes sense: we expect the bar to funnel gas towards the central regions of the galaxy (e.g. Kim et al., 2024; Verwilghen et al., 2024), and often observe higher molecular gas surface densities in central regions as a result (e.g. Fisher et al., 2012; Yu et al., 2022). Through the bar structure and associated torques, gas loses angular momentum, allowing it to sink to the centre of the galaxy, creating star-forming structures including rings, filled discs, or nuclear spirals that in turn build a nuclear disc (e.g. Seo et al., 2019; Lin et al., 2020).

### v2

@@ 3.3.3 Skew, $h_{3}$, maps and profiles
3.3.4 Kurtosis, $h_{4}$, maps and profiles
In Figures 314, we present the line-of-sight kurtosis, or $h_{4}$, maps. There are fewer spatially coherent features visible in these maps, possibly the result of observational challenges related to retrieving high-order velocity moments from medium-resolution spectra (e.g. Wang et al., 2024). Nevertheless, we do notice some trends. Given their co-spatial nature, we expect that the large longitudinally-extended regions of high positive $h_{4}$ around the centres of some galaxies that correspond to low $\sigma_{\star}$ and clear deviations from symmetry in the surface brightness contours are caused by the dust lanes. In some galaxies, including PGC 044931, NGC 3957, NGC 0522, and ESO 079-003, we see regions of high positive $h_{4}$ extending from the central regions of the galaxy 1–2 kpc off the z=0 line. In PGC 044931, IC 1711, and NGC 3957, the small, inner areas of opposing $h_{3}$ and $V_{\star}$ sign are co-spatial with a coherent region of high positive $h_{4}$. The strength of these features with respect the the surrounding disc are highlighted in the 1D profiles of Figure 16b.
@@ 4 Discussion
Here we attempt to connect the observations of Section 3.3 to an understanding of the central kinematic sub-structures in the GECKOS galaxies. Already clear is the diversity of structures in the $V_{\star}$ maps (visible central kinematically-decoupled rapidly-rotating structures), $\sigma_{\star}$ maps (structure and intensity of sigma peaks at the centre of the galaxy, effects of dust lanes, central ‘croissant’ features), $h_{3}$ (coherent areas of matching/opposing sign with $V_{\star}$), and $h_{4}$ (small, highly-peaked central regions). That said, some features are similar throughout the sample: all galaxies possess a rotating disc structure and some degree of $h_{3}$ and $V_{\star}$ sign mismatch in the outer disc regions of the galaxy. In the following, we explore these observations as possible tracers of kinematic sub-structures.
@@ 4.1.1 Midplane features
  • A ‘double-hump’ rotation curve, with a stellar mean line-of-sight velocity ($V_{\star}$) profile exhibiting an initial rise, followed by a hump or plateau, and then a secondary rise, before flattening out. The authors surmise that the hump is caused by (mostly inner) bar $x_{1}$ orbits, the elongated family of orbits that follow the major axis of the bar (e.g. Contopoulos & Papayannopoulos, 1980; Athanassoula, 1992), adding a distinct component to the LOSVD.
  • A flat-topped or weakly peaked radial velocity dispersion ($\sigma_{\star}$) profile, sometimes with a local central minimum. Broad ‘shoulders’ or secondary maxima may be present in strong bars, usually towards the end of the bar, just inside the inner ring. Bureau & Athanassoula (2005) attributed the central $\sigma_{\star}$ peak to the large variety of orbital shapes encountered in strongly-barred discs, particularly along the major axis. They posited that the $\sigma_{\star}$ ‘shoulders’ are produced by the tips of the $x_{1}$ orbits and the inner 4:1 orbits. In particular, $x_{1}$ orbits with loops around the major axis are expected in strong bars, causing a local increase in the velocity dispersion, strongest for bars seen end-on (Athanassoula, 1992). The higher energy inner 4:1 orbits have similar loops, which increase $\sigma_{\star}$ for bars seen both side-on and end-on.
@@
Crucially, Bureau & Athanassoula (2005) also show that the strength of these kinematic features is correlated with bar viewing angle, such that the strongest signals are seen when the bar is oriented end-on to the viewer. Through observations of 30 edge-on spiral galaxies, Chung & Bureau (2004) provided an alternative explanation for the ‘double-hump’ rotation curve by taking into account the central minimum behaviour of the $\sigma_{\star}$ profile and $h_{3}$-$V_{\star}$ profile anti-correlation in this region. From these observations, they suggest that the inner regions of the galaxy are both cold, and also very close to axisymmetric. They speculate that these regions have largely decoupled from the rest of the galaxy (and the bar) and circularised, to form a nuclear disc.
@@
Comparing the $h_{3}$ 1D profiles of Figure 16a with $V_{\star}$ in Figure 15a, some $h_{3}-V_{\star}$ correlations are seen in the central few kpc of PGC 044931, NGC 3957, IC 1711, IC 5244, ESO 484-036, and ESO 079-003. Taken with the trends seen in the $V_{\star}$ and $\sigma_{\star}$ profiles, we can reliably report that these galaxies host strong evidence of bar structure from their stellar kinematics along the z=0 line.
Finally, we discuss the effects of dust on the 1D line profiles. Prominent dust lanes are seen in every galaxy in this sample. When this dust absorbs or scatters stellar photons, it does the same to kinematical information contained within these photons. Baes et al. (2003) present model galaxy rotation curves and $\sigma_{\star}$ profiles both with and without the application of radiative transfer models incorporating absorption and scattering by interstellar dust grains. They find that for highly inclined galaxies, rotation curves and central $\sigma_{\star}$ peaks are significantly flattened, leading to significant underestimation of the true rotational velocity of the inner regions (e.g. Davies, 1990; Bosma et al., 1992; Catinella et al., 2006). For this reason, we do not further quantify the shapes of the rotation curves observed beyond the presence of ‘double hump’ features, but do note that Baes et al. (2003) report that even a slight deviation from exactly edge-on inclinations dramatically reduces dust effects.
@@ 4.1.2 Off-plane features
Iannuzzi & Athanassoula (2015) extended previous long-slit work by providing 2D, projected line-of-sight kinematics for a set of barred/BP bulge dynamical models featuring star formation, one of which also hosted a spherical bulge. The velocity maps presented in that work, whilst appearing qualitatively similar to those seen here, differ in the uniform nature of the kinematic minor axis. In all cases, the kinematic minor axis was aligned perpendicular to the photometric major axis. In the GECKOS galaxies, we often see central ‘wedge-shaped’ deviations, as illustrated in e.g. NGC 0522. These deviations are generally coincident with the dust lane, as shown by the overlaid flux contours on the $\sigma_{\star}$ maps, and so it is reasonable to attribute them to dust. The kinematic minor axis is straight in the model predictions as regardless of bar viewing angle, all star particles on bar orbits will be accounted for, regardless of their position on the near or the far side of the disc. Dust breaks the bi-symmetry of a bar (and indeed any disc), as light from stars on the near side of the bar will contribute far more than those from the far side obscured by dust. This net radial flow manifests itself as the ‘wedge’ shape of the kinematic axis of GECKOS galaxies PGC 044931, NGC 0522, ESO 079-003, all of which are barred, but also NGC 0360 which we have not classified as barred. Further analysis of these structures in conjunction with complementary ALMA data will give us additional information on the location of the near side of the bar, and is planned for future work.
Relatedly, we examine the lighter-coloured regions of the $V_{\star}$ maps, indicative of a velocity magnitude lower than the surroundings. These regions are co-spatial with the dust lane, as seen from the surface brightness contours overlaid on the $\sigma_{\star}$ maps. As illustrated in Davies (1990), the component of rotational velocity recovered from a highly inclined galaxy with an optically thick inner region is lower than expected. Coupled with observational results showing that dust-free rotation curve measures such as HI and CO generally produce steeper inner rotation curves compared to optical measures (e.g. Bosma et al., 1992), we assume that the regions of lower velocity magnitude are the direct result of dust attenuation.
Finally, we mention also the off-plane deviations from a straight $V_{\star}=0$ isovelocity line in PGC 044931, NGC 0360, ESO 120-016, NGC 5775, ESO 079-003, and ESO 484-036. These deviations are clearest on the lower sides of the velocity maps, given that we calculated the systemic velocities from the top half. Given that the lower half of the $V_{\star}$ maps contains the dust lane (if the galaxy is not perfectly edge-on), we expect this side to be more dust-obscured than the top side. We hypothesise that it is the greater degree of dust obscuration that causes these deviations from a perfectly straight $V_{\star}=0$ isovelocity line.
@@
In addition to confirming 1D predictions and observations, Iannuzzi & Athanassoula (2015) found that the 2D line-of-sight mean velocity maps do not provide a robust predictor of the presence of BP bulges as the degree of cylindrical rotation (i.e. mean stellar velocity independent of height above the disc) varies. In their work, Iannuzzi & Athanassoula (2015) reported a tendency for the degree of cylindrical rotation to weaken when simulations are seen from side-on to progressively more end-on bar angles, though the importance of this effect varies from case to case without a clear trend in BP bulge strength. A variation in BP bulge rotational properties is also reported by Athanassoula & Misiriotis (2002), who attribute this behaviour to different disc and halo contributions to the total circular velocity curve in their various models. Williams et al. (2011)’s observations support this idea with observations of five edge-on discs hosting BPs that show a variety of projected off-plane velocity behaviours. However, Molaeinezhad et al. (2016) developed a new technique to measure the degree of cylindrical rotation and found a wide range of cylindrical rotation, but on average, BP bulges displayed higher values for this parameter than non-BP bulges. We do not look further into cylindrical rotation indicators in this work, though note that the GECKOS sample will provide an excellent opportunity to understand the off-plane velocity behaviour of BP bulges.
@@
  • Elongated ‘wings’ of large values $h_{3}$. Again, the bar viewing angle is important for both the amplitude and morphology of $h_{3}$ maps. Iannuzzi & Athanassoula (2015) and Li et al. (2018) analyse separate N-body models of barred systems hosting a BP bulge at several different bar viewing angles and conclude that while bar-driven signatures in $h_{3}$ are barely visible when the bar is side-on to the observer, considerable changes occur as the bar is rotated, such that large regions of $h_{3}$ maxima and minima develop roughly in the regions corresponding to the peanut, increasing in size such that they are largest at intermediate bar viewing angles (30-60°), and magnitude such that they are strongest when the bar is viewed end-on.
  • An ‘X’-shaped region of deep $h_{4}$ minima, becoming deeper and better spatially defined as bar viewing angle moves from side-on to end-on.
Both $h_{3}$ and $h_{4}$ features become stronger at larger BP bulge strengths (for a given bar viewing angle) in edge-on galaxies.
Taking these predictions to the GECKOS sample, we see coherent areas of $h_{3}$ of same sign as $V_{\star}$ at the centres of all BP bulge galaxies (though to varying degrees). We see an excellent example of the predicted wing structure in IC 5244 and IC 1711. These areas of coherent $h_{3}$ sign originate in the central galactic regions and extend up to 2 kpc off the disc, and up to 5 kpc radially. We expect that these regions correspond to the area of influence of the bar, with the radial width related to the bar viewing angle.
@@
Iannuzzi & Athanassoula (2015) applied Voronoi binning to their particle data before fitting the LOSVD with a Gauss-Hermite to obtain their kinematics; an equivalent strategy to Wang et al. (2024). We believe that the reason that these simulation studies found strong $h_{4}$ features that are not seen in the GECKOS data is most likely due to observational limitations, i.e., the process of transferring particle statistics to spectra, and then extracting the $h_{4}$ from weak spectral features. Due to this process, all of the ‘real’ $h_{4}$ features become more shallow or disappear. We refer the reader to Sections 3.3 and 4.1 of Wang et al. (2024) for a more thorough discussion on this point.
4.1.3 Inclination effects
Iannuzzi & Athanassoula (2015) investigate the effects of galaxy inclination on high order kinematic maps. When a galaxy is inclined to be perfectly edge-on to the viewer, a radially-extended ($2\lesssim\rm{R[kpc]}\lesssim 7$ ) region of $h_{3}$ maxima corresponding to a sign match with $V_{\star}$ occurs along the z=0 line. In the 80° case, these regions move off the z=0 line to $\sim$2 kpc, extending upwards of 5 kpc both above and below the z=0 line. The $h_{3}$ maxima along the z=0 line changes sign, though remains of high magnitude.
Bureau & Athanassoula (2005) report similar behaviour along the midplane, attributing the switch in sign from a $h_{3}$-$V_{\star}$ correlation to an anti-correlation to the projected signature of the outer disc disappearing, which in turn increases the observed mean velocities, and shifts the asymmetric tail of material in the LOSVDs from high to low velocities. We observe qualitatively similar behaviour to the Iannuzzi & Athanassoula (2015) 80° inclination case in NGC 3957, NGC 0522, PGC 044931, IC 5244, and ESO 079-003, all of which exhibit $h_{3}$ maxima corresponding to a sign match with $V_{\star}$ located as coherent regions both above and below the z=0 line. Given the slight inclinations apparent from the unsharp masked optical images and Spitzer maps, we expect that the majority of the GECKOS galaxies possess inclinations between 80–90°. Work to accurately characterise the GECKOS sample inclinations via detailed orbital modelling techniques are underway, however we speculate that a slight deviation from perfectly edge-on geometry is the cause of these regions of $h_{3}$-$V_{\star}$ sign match moving off the z=0 line.
@@ 4.1.4 Bar viewing angle
$h_{3}$ maps vary considerably in both amplitude and morphology with bar viewing angle, according to both the Iannuzzi & Athanassoula (2015) and Li et al. (2018) model predictions. The physical extent and maximum value of the observed $h_{3}$ ‘blobs’ appears least conspicuous in the side-on case, and increases in both size and magnitude (whilst still maintaining the appearance of extended blobs), as the bar is rotated toward the viewer. The $h_{3}$ blobs are most prominent when the bar is seen end-on. In Figures 3-14, we see the strongest degree of coherent $h_{3}$ structure (correlating in sign with regions of $V_{\star}$) in the central regions of IC 5244 and IC 1711, adding weight to the idea that these galaxies possess end-on bars. We see the lowest degree of $h_{3}$ structure in ESO 120-016 and NGC 0522. While it is possible that ESO 120-016 is indeed not a BP bulge galaxy, NGC 0522 possesses a definite X shape in the DECaLS and Spitzer imaging. For this reason, we expect that NGC 0522 must possess a bar that is close to side-on orientation. For all other galaxies, we infer from the $h_{3}$ maps that the bar viewing angle is somewhat intermediate between side-on and end-on.
@@ 4.2 Evidence for nuclear discs
Following from the predictions of Bureau & Athanassoula (2005), an additional predictor of bar structure along the kinematic major axis of edge-on galaxies was provided by Chung & Bureau (2004), who illustrate the features mentioned in Section 4.1 in their Figure 9, and observed a strong anti-correlation between $h_{3}$ and $V_{\star}$ profiles in the innermost region of BP bulge galaxies, which they attributed to the presence of a cold, rapidly-rotating nuclear disc built by a bar. This anti-correlation switches back to a correlation outside of the nuclear disc region but still within the bar region, in line with the Bureau & Athanassoula (2005) prediction.
@@
Evidence for the presence of nuclear discs is observed in the $V_{\star}$, $\sigma_{\star}$, $h_{3}$, and $h_{4}$ maps of IC 1711, IC 5244, NGC 3957, and PGC 044931. Whilst all show modest increases in rotational velocity in small regions at their very centres, NGC 3957 and PGC 044931 also show a ‘croissant’-shaped feature in their velocity dispersion maps, with a central depression of lower $\sigma_{\star}$. All four galaxies also show $h_{3}$-$V_{\star}$ sign mismatches in these central regions, co-located with regions of high positive $h_{4}$.
@@
These features are perhaps better shown in the 1D line profiles of Figures 15 and  16, where the four galaxies in question lie at the top of these figures. The central minima in the dispersion profiles of PGC 044931, NGC 3957, and IC 1711 are features that are sometimes referred to as ‘sigma drops’ (e.g. Emsellem et al., 2001). Observationally, this feature has been frequently seen in long-slit spectra (e.g. Emsellem et al., 2001; Márquez et al., 2003; Chung & Bureau, 2004; Comerón et al., 2008; Méndez-Abreu et al., 2014) and IFS data (e.g. Crocker et al., 2011; Lin et al., 2017; Pinna et al., 2019b; Shimizu et al., 2019). We expect that these $\sigma$ drops are due to the presence of a dynamically cold and kinematically-decoupled central stellar disc originating from gas inflow. Interestingly, upon closer inspection of the $\sigma_{\star}$ radial profiles of PGC 044931, NGC 3957, and IC 1711, we see that their profiles bifurcate in the central regions, with one population of stars of low $\sigma_{\star}$ (the nuclear disc stars), and one population with higher $\sigma_{\star}$. In fact, in IC 5244 we see only a $\sigma_{\star}$ peak, with perhaps only a tiny hint of a drop. We expect that the peaks are due to bar orbits, as discussed in Section 4.1.
@@
The N-body models of Bureau & Athanassoula (2005) do not include nuclear discs, but do reproduce this central $\sigma_{\star}$ minima, attributing it to the orbital structure of strongly barred discs. More recent N-body and SPH simulations that include stars, gas, and star formation show that young stars born in the nuclear regions from dynamically cold gas have a velocity dispersion lower than the older stellar population (Wozniak et al., 2003; Michel-Dansac & Wozniak, 2004; Wozniak, 2007; Cole et al., 2014; Portaluri et al., 2017), and hence attribute the sigma drop feature to a dynamically cold, nuclear disc. Contextually, the presence of a nuclear disc in a barred galaxy makes sense: we expect the bar to funnel gas towards the central regions of the galaxy (e.g. Kim et al., 2024; Verwilghen et al., 2024), and often observe higher molecular gas surface densities in central regions as a result (e.g. Fisher et al., 2013; Yu et al., 2022). Through the bar structure and associated torques, gas loses angular momentum, allowing it to sink to the centre of the galaxy, creating star-forming structures including rings, filled discs, or nuclear spirals that in turn build a nuclear disc (e.g. Seo et al., 2019; Lin et al., 2020).
The $h_{3}$ and $h_{4}$ maps also encode information on the presence of nuclear discs at the centres of some of the GECKOS BP bulge galaxies. There is a small region of high positive and negative $h_{3}$ at the centres of IC 5244, IC 1711, NGC 3957, and PGC 044931, all of which correspond to an opposing sign in the same regions of the $V_{\star}$ maps. Presumably, this feature is related to the central $h_{3}$-$V_{\star}$ anti-correlation seen by Chung & Bureau (2004), who explain this observation via a region that appears to have largely decoupled from the rest of the galaxy (and the bar) and circularized, forming a dense and (quasi-) axisymmetric central stellar disc. This feature was also observed in most of the TIMER galaxies (Gadotti et al., 2020), NGC 1381 (Pinna et al., 2019b), NGC 5746 (Martig et al., 2021), and in NGC 4643 (Erwin et al., 2021). We also observe coherent regions of $h_{4}$ maxima co-spatial with the central $h_{3}$-$V_{\star}$ sign mismatch and $\sigma$ drop regions. We expect that we are observing the projected equivalent of the regions of $h_{4}$ maxima seen in most TIMER survey galaxies Gadotti et al. (2020) (also e.g. NGC 4643 Erwin et al., 2021). These are explained by Gadotti et al. (2020) as areas in which there is a superposition of multiple kinematic components - in this case, a low-dispersion nuclear disc component on top of a relatively higher-sigma underlying disc component.